OpenStudy (anonymous):

HELP PLEASE!!!! If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. -x + 5y = 1 2x + 4y = -4

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

hey can you help me

terenzreignz (terenzreignz):

Okay, substitution, the most basic method, possibly the easiest, but not always the most efficient :) You have two equations -x + 5y = 1 and 2x + 4y = -4 The way to do this is to solve for x in the first equation (or second, really, but first is easier). What I mean is that rearrange the first equation so that x stands alone in one side of the equal sign.

OpenStudy (anonymous):

a. 2(5y - 1) + 4y = -4 b. 2(-5y + 1) + 4y = -4 c. 2x + 4(-5y + 1) = -4 d. 2x + 4(5y - 1) = -4 these are the possible answers

terenzreignz (terenzreignz):

Yeah... might be to your benefit to pretend we don't have choices for now. Let's look at the first equation. -x + 5y = 1 If we subtract 5y from both sides, what do we get?

OpenStudy (anonymous):

1-5y=-x

terenzreignz (terenzreignz):

That's right. But that's a -x How do we get rid of that negative sign?

OpenStudy (anonymous):

multiply both sides by neg1

terenzreignz (terenzreignz):

and you get...?

OpenStudy (anonymous):

2x=-1+5y

terenzreignz (terenzreignz):

Why 2x? How did you get 2x? -x = 1 - 5y multiply both sides by -1 -1(-x) = -1(1 - 5y)

OpenStudy (anonymous):

ok so x=-1+5y

terenzreignz (terenzreignz):

or x = 5y - 1 Now, given your SECOND equation 2x + 4y = -1 replace the x here with the right-hand side of your first equation modified: x = 5y - 1

OpenStudy (anonymous):

2(5y-1)+4y=-1

terenzreignz (terenzreignz):

Sorry, it was 2x + 4y = -4

OpenStudy (anonymous):

thats isn't even one of the answers

terenzreignz (terenzreignz):

I know. I made a typo. 2x + 4y = -4 and replace the x with the expression in terms of y. x = 5y - 1

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

abc or d

OpenStudy (anonymous):

so d

terenzreignz (terenzreignz):

You replace the x in 2x + 4y = -4 with 5y - 1

OpenStudy (anonymous):

so a

terenzreignz (terenzreignz):

Yeap...

OpenStudy (anonymous):

hey i just posted another ? can you come help me